Find all real solutions of the equation.
x^3+6x^2-3x-18=0

Answer by edjones(8007)

x^3+6x^2-3x-18=0
x^2(x+6)-3(x+6)=0
(x^2-3)(x+6)=0
x^2=3
x=sqrt(3), x=-sqrt(3), x=-6
.
Ed